jules needs to purchase a ladder that reaches the top of a 16-foot building. if the bottom of the ladder will be placed 12 feet from the base of the building, how long should jules's ladder be?
Answer by richwmiller(17219)

(3*4)^2+(4*4)^2=(5*4)^2
144+256=400
This is a classic 3,4 5 right triangle multiplied by 4